# Diffusion-Aided Joint Source Channel Coding For High Realism Wireless Image Transmission

### v2

@@ II-C1 Preliminary of Latent Diffusion Model
Ho et al. introduced the Denoising Diffusion Probabilistic Model (DDPM), a generative approach that iteratively refines a sample from a noise distribution towards a target data distribution using a sequence of denoising operations [28]. To improve this process for efficiency and training stability, Rombach et al. conceived the Stable Diffusion method wherein a variational autoencoder (VAE) initially condenses images into a lower-dimensional latent space, and subsequent diffusion steps occur [39]. The forward diffusion process employs a Markov chain that incrementally introduces Gaussian noise to the latents at each time step $t$, with a variance denoted by $\beta_{t}\in(0,1)$, applied to the state $z_{t-1}$ from the previous step. Summarized over the time steps from $1$ to $t$, this process is represented as:
@@
Here, $z_{0}=E_{ldm}(x)$ is the latent representation of the original image encoded by the VAE, $\epsilon$ is sampled from a standard Gaussian distribution $\mathcal{N}(0,\mathbf{I})$, $\alpha_{t}=1-\beta_{t}$, and $\bar{\alpha}_{t}$ is the cumulative product of $\alpha_{i}$ from the initial time step to time step $t$, reflecting the progression of the diffusion process.
The aim of the denoising process is to estimate the prior latent state $z_{t-1}$ from the current state $z_{t}$, guided by the latent’s original state $z_{0}$. The posterior distribution for this estimation, $p(z_{t-1}|z_{t},z_{0})$, is modeled as a Gaussian $\mathcal{N}(z_{t-1};\mu_{t}(z_{t},z_{0}),\sigma_{t}^{2}\mathbf{I})$. The mean of this distribution, $\mu_{t}(z_{t},z_{0})$, can be parameterized as:
@@
Sampling $z_{t}$ hinges on accurately approximating the noise component $\epsilon$, which is facilitated by a neural network denoted as $\epsilon_{\theta}$. The network is typically conditioned on a context variable $c$ that may embody supplementary information such as text descriptions. Consequently, the network $\epsilon_{\theta}$ is trained to minimize the discrepancy between the true noise $\epsilon$ and its estimated value, using the loss function:
@@ II-C2 The Denoiser Structure
The proposed denoiser $\epsilon_{\theta}$ is illusttrated in Figure 3. It utilizes the UNet architecture [44] from the Stable Diffusion denoiser (shown in blue) which remains unchanged during training. This architecture comprises of four encoder blocks, four decoder blocks, and one middle block. Each encoder/decoder block includes two main blocks and one convolutional block for scaling operations. The middle block contains two main blocks. Each main block consists of 4 ResNet and 2 Vision Transformers (ViTs). The ViTs integrate self-attention and cross-attention mechanisms to facilitate interaction with external prompts, such as text encoded by the CLIP text encoder.
@@
The control module highlighted in green in Figure 3 is a pivotal component of the proposed denoiser architecture. Inspired by ControlNet [37] and LAControlNet[40], this module adopts the UNet encoder architecture and middle block used in the Stable Diffusion model. The control module is comprised of four encoder blocks in addition to the middle block to handle four distinct resolutions—64×$\times 64$ 64, $32\times 32$, $16\times 16$, and $8\times 8$—capturing a wide range of feature abstraction scales. The outputs from these blocks are fed into the four decoder blocks and the middle block of the Stable Diffusion Unet. Before being fed into the UNet, the control module’s outputs go through a series of $1\times 1$ convolutional layers (‘zero conv’ layers in Figure 3).
@@
It is worth noting that we use the pre-trained Stable Diffusion Unet (in blue in Figure 3) in the denoiser without retraining (fine-tuning) whereas we train the control module (in green) for our proposed control module input. We obtain the control module input by concatenating the latent representation $z_{t}$ with the spatial features $f_{spa}$ from the LDM encoder $E_{ldm}$. This concatenation enhances the input’s representational capacity but it extends the control module network’s channel depth necessitating additional parameters beyond those present in the pre-trained model. To handle this extension, the extra parameters are set to zero at initialization to maintain a neutral effect at the start of the training process.
To encode the channel condition SNR denoted by $\gamma$, we use a specialized embedding transformation to generate SNR embeddings that have the same dimensionally as the time embedding, therefore ensuring seamless integration with the control module. Furthermore, textual features $f_{txt}$ serve as crucial prompts infusing contextual information to both the pre-trained UNet and the control module. As shown in Figure 3, textual features $f_{txt}$ are used as text prompts and are passed to the ViTs within each encoder/middle/decoder block as queries. Similar to the LDM loss (11), our approach focuses on reducing the Mean Square Error (MSE) of the noise prediction while adhering to the condition set $\bar{c}$ given in equation (8). Our loss function is thus represented by:
@@
  Input: Time steps $T$, spatial features $f_{spa}$, textual features $f_{txt}$, channel SNR $\gamma$, scale factor $\lambda$, denoiser $\epsilon_{\theta}$
@@
       $\hat{z}_{0}\leftarrow\tilde{z}_{0}-\frac{\lambda}{C_{l}H_{l}W_{l}}(\tilde{z}_{0}-f_{spa})$
@@ II-C3 Adding Intermediate Guidance
Diffusion models, as articulated by Lin et al. [40], strive to carefully navigate the delicate trade-off between the realism of the generated images and their fidelity to the original content. It is worth mentioning that these models might unintentionally introduce mismatched elements that may look realistic but different from the actual target. This issue is more prevalent especially when there is a significant difference between the target domain and the training domain. To address this issue and ensure that the reconstructed images faithfully reflect the original content, we adopt an optional guidance mechanism similar to [40] that leverages the initial JSCC reconstruction $\hat{x}$ to steer the intermediate latent state $\tilde{z}_{0}$ during the denoising process. This guidance allows for the computation of $\tilde{z}_{0}$ in equation (9) as the intermediate state estimation at a particular time $t$.
@@
$\hat{z}_{0}=\tilde{z}_{0}-\frac{\lambda}{C_{l}H_{l}W_{l}}(\tilde{z}_{0}-f_{spa}),$ (14)
where the variables $C_{l}$, $H_{l}$, and $W_{l}$ correspond to the dimensions of the latent space, and $\lambda$ is a scalar that regulates the level of alignment between $\hat{z}_{0}$ and $f_{spa}$. The posterior distribution $p(z_{t-1}|z_{t},\hat{z}_{0})$ subsequently determines the sampled state $z_{t-1}$ as given by equation (10) substituting $z_{0}$ with $\hat{z}_{0}$.
This entire process is captured in Algorithm 1. It is worth emphasizing that the application of the intermediate guidance is optional and can be omitted by assigning a value of zero to $\lambda$, thereby maintaining the general functionality of the diffusion process without the additional guidance.
@@ III Training and Evaluation
Figure 4: Evaluation of the proposed approach versus baselines across different transmission rates under 1dB (top) and 10dB SNR (bottom) on the Kodak dataset.
Figure 5: Evaluation of the proposed approach versus baselines across different SNR levels with transmission rates of $1/128$ cpp (top) and $1/32$ cpp (bottom) using the Kodak dataset.
Figure 6: Visual comparison between DiffJSCC and baselines on the Kodak dataset at 1dB SNR. For DiffJSCC, we show both the final reconstructed images and the captions generated from the initial reconstruction.
@@ III-A1 Dataset
Our DiffJSCC framework evaluation encompasses two principal applications: general image transmission and specialized facial image transmission. In the general image transmission scenario, DiffJSCC is fine-tuned on a selected dataset of approximately 300,000 high-resolution images from the OpenImages collection [^1] https://storage.googleapis.com/openimages/web/index.html. Throughout the training phase, these images are subject to random cropping operations, resulting in dimensions of $256\times 256$ pixels. For the performance evaluation, we use the standard Kodak dataset, which consists of 24 curated high-fidelity images with dimensions of $768\times 512$ pixels. In addition, we extend the performance evaluation to include the ADE20K validation set with 2000 images to assess image reconstruction quality and semantic segmentation accuracy.
@@ III-A2 Implementation
Within our framework, we utilize an SNR-adaptive JSCC architecture for both the encoder and decoder, as detailed by Yang et al.[22]. This architecture incorporates the SNR channel state information into the feature processing pipeline via a Squeeze-Excitation mechanism [21], enhancing the adaptability to varying SNR conditions. The transmission rate $R$ is fixed for each model configuration without the policy network introduced in [22]. Each model has a fixed rate $R$ which is determined by the output channel count ($C_{out}$) and the downsampling factor ($D$), satisfying $R=C_{out}/512$ as $D$ is set to 4. General image transmission training involves random $256\times 256$ patches and a batch size of 32, while the facial image dataset utilizes full $512\times 512$ CelebAHQ images with a smaller batch size of 16. An Adam optimizer initializes learning at $10^{-3}$, tapering over 100,000 iterations following a cosine decay schedule.
@@
We used 2 NVIDIA A40 GPUs to accelerate training. Though the Stable Diffusion model is trained using 1000 time steps, we use a spaced DDPM sampling for 50 steps during inference, a common strategy [39, 40] to improve sampling efficiency. Unless specified otherwise, the scaling factor $\lambda$ is preset to zero as a standard baseline configuration without the intermediate guidance introduced in Section II-C3.
@@ III-B Metrics
For evaluations, we adopt standard distortion metrics such as PSNR and MS-SSIM[45] to measure image fidelity concerning pixel intensity and structural details, respectively. To evaluate the perceptual quality of reconstructed images, we employ more recently proposed metrics such as Learned Perceptual Image Patch Similarity (LPIPS)[46] and Fréchet Inception Distance (FID) scores[47]. These metrics provide crucial insights into visual similarity and how well the distribution of generated images aligns with that of real images. Given the Kodak dataset contains only 24 images, we transmit each image five times to reliably measure our performance metrics. For Kodak images, the FID score is calculated over half-overlapping patches of $256\times 256$ pixels. Our evaluation also includes the performance of the semantic segmentation task on the generated and original images measured by the mean Intersection over Union (mIoU) metric.
@@ III-C Performance on General Image Transmission
DeepJSCC and DiffJSCC methods support four distinct rates: $R=1/128,1/64,1/32,$ and $1/16$. Notably, the basic DeepJSCC method surpasses traditional separate coding methods across multiple metrics, validating previous studies. The introduction of generative diffusion-based refinements from our DiffJSCC may induce degradation in traditional distortion metrics such as PSNR and MS-SSIM. Nevertheless, DiffJSCC exhibits remarkable improvements in visual quality assessed by LPIPS and FID metrics. These metrics capture the capability of DiffJSCC to produce better photorealistic reconstructions at compression levels unachievable by the other schemes (including theoretical ‘BPG-Capacity’).
@@
Notice, at a 1dB SNR, DiffJSCC attains a low FID score of $117.9$ at the transmission rate of merely $R=1/128=0.0078$, outperforming the ‘BPG+Capacity’ baseline’s FID score of $165.4$ at a substantially higher rate of $R=0.113$. This translates to DiffJSCC achieving superior image quality by a margin of $28.7\%$ while utilizing 14×14$\times$14 × less symbols. In scenarios where the SNR is increased to 10dB, DiffJSCC continues to hold a significant advantage over the other methods although the performance gap in terms of LPIPS and FID is reduced. When transmitting at a rate of $R=0.0078$, the FID score of DiffJSCC demonstrates an impressive $64.3\%$ enhancement over DeepJSCC and a $70.7\%$ over ‘BPG+Capacity’. This confirms the effectiveness of the DiffJSCC framework in reconstructing highly realistic images under stringent rate and channel constraints.
@@
In Figure 5, we assess DiffJSCC in comparison to other deep JSCC approaches across various SNR conditions. The DeepJSCC-$adv$ model from Yang et al.[13] is included for comparison, which uses the hyperparameter $\alpha$ to adjust the weight of adversarial loss. Various $\alpha$ parameters were tested to find the optimal trade-off between distortion and visual quality. Similarly, we experimented with different $\lambda$ values in DiffJSCC with the same goal. These evaluations were conducted at transmission rates of $R=1/128$ and $R=1/32$, with SNR values set at $\gamma=1,4,7,10,13$ dB. All benchmarks including ours are based on the same deep JSCC network architecture. The results in Figure 5 demonstrate that although the DeepJSCC-$adv$ baseline successfully lowers LPIPS, it struggles to effectively minimize FID scores regardless of the $\alpha$ value used. The FID score remained above $190$ at $R=1/128$ and exceeded $110$ at $R=1/32$. Conversely, DiffJSCC significantly improves both LPIPS and FID scores over the entire range of SNR values, while maintaining similar PSNR and MS-SSIM distortion metrics. At $\lambda=0$ and $R=1/128$, DiffJSCC shows a notable average improvement of $41.4\%$ in LPIPS and $56.7\%$ in FID scores compared to the standard DeepJSCC. When compared with DeepJSCC-$adv$ using $\alpha=5\times 10^{-3}$, it achieves $21.1\%$ better LPIPS and $52.2\%$ enhanced FID metrics. At $R=1/32$, DiffJSCC consistently demonstrates a significant gain in performance, improving average LPIPS and FID scores by $43\%$ and $53.2\%$ over DeepJSCC, respectively, and $18.9\%$ and $48.4\%$ compared to DeepJSCC-$adv$. Increasing the scale factor $\lambda$ from 0 to 100 gradually aligns the generated image with the initial JSCC reconstruction, enhancing PSNR and MS-SSIM but reducing perceptual quality. However, this trade-off still leads to better visual perception metrics (LPIPS and FID) than the baselines. This significant improvement in the perceived visual quality highlights the effectiveness of our fine-tuning control module along with the Stable Diffusion model.
Figure 7: Results on the ADE20k dataset with $R=1/128$ and $R=1/32$. Higher mIoU scores indicate better segmentation.
Figure 8: Visualizations on ADK20K with $R=1/128$ and 1dB SNR.
@@
Figure 6 provides visual examples of the superior perceptual quality achieved by DiffJSCC at a challenging SNR of 1dB. At a transmission rate of $R=1/128$, the reconstructions produced by the conventional DeepJSCC framework exhibit significant distortion and have difficulty recovering important image details. Conversely, DiffJSCC demonstrates a remarkable capacity to restore crucial elements of the image, like the structures of the lighthouse and house, offering significantly enhanced perceptual quality despite slight variances from the original. With the transmission rate raised to $1/32$, DiffJSCC’s capabilities become more evident, resulting in reconstructions that accurately reflect the original images. Furthermore, we demonstrate the validity of captions generated by the Blipv2 model based on the initial JSCC reconstructions. These captions prove to be robust against variations in reconstruction quality, consistently capturing the important content keywords of each image (highlighted in red). This confirms the ability of DiffJSCC to convey essential semantic information even at reduced transmission rates.
@@
Figure 9: Comparison between the proposed method and baselines at various SNRs on CelebAHQ.
@@
Figure 10: Visualizations on CelebAHQ with $R=1/256$ in both high and low SNR scenarios.

### v3

@@ II-B Multimodal Condition Generation
Figure 3: Network structure of the proposed denoiser $\epsilon_{\theta}$ in the conditional latent diffusion model. The pre-trained Stable Diffusion model is shown in blue and the fine-tuned control module is shown in green.
@@ II-C Conditional Latent Diffusion Model
The proposed DiffJSCC framework takes the set of conditions $\bar{c}$ as the input of the conditional latent diffusion process as depicted in Fig. 2. The diffusion process starts with a random input $z_{T}$ sampled from $\mathcal{N}(0,\mathbf{I})$ and iterates the denoising process with the multi-modal condition set $\bar{c}$ using the proposed denoiser (Fig. 3) until the denoised output $z_{0}$ is obtained. The final output image $\hat{x}_{diff}$ is generated using an LDM decoder on $z_{0}$: $\hat{x}_{diff}=D_{ldm}(z_{0})$.
@@ II-C1 Preliminary of Latent Diffusion Model
Ho et al. introduced the Denoising Diffusion Probabilistic Model (DDPM), a generative approach that iteratively refines a sample from a noise distribution towards a target data distribution using a sequence of denoising operations [29]. To improve this process for efficiency and training stability, Rombach et al. proposed the Stable Diffusion method in which a variational autoencoder (VAE) initially maps images into a lower-dimensional latent space, followed by diffusion denoising steps [43]. The forward diffusion process employs a Markov chain that incrementally introduces Gaussian noise to the latents at each time step $t$, with a variance denoted by $\beta_{t}\in(0,1)$, applied to the state $z_{t-1}$ from the previous step, which is:
$z_{t}=\sqrt{1-\beta_{t}}z_{t-1}+\sqrt{\beta_{t}}\epsilon,$ (10)
where $\epsilon$ denotes the noise sampled from a standard Gaussian distribution $\mathcal{N}(0,\mathbf{I})$. Summarized over the time steps from $1$ to $t$, the forward process is represented as:
@@
Here, $z_{0}=E_{ldm}(x)$ is the latent representation of the original image encoded by the VAE, $\alpha_{t}=1-\beta_{t}$, and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ is the cumulative product of $\alpha_{i}$ from the initial time step to time step $t$, reflecting the progression of the diffusion process.
The aim of the denoising process is to estimate the prior latent state $z_{t-1}$ from the current state $z_{t}$. According to [29], the probability $p_{\theta}(z_{t-1}|z_{t})$ can be solved by minimizing the Evidence Lower Bound (ELBO), which is equivalent to minimizing the KL-divergence between $p_{\theta}(z_{t-1}|z_{t})$ and the posterior distribution $q(z_{t-1}|z_{t},z_{0})$. $q(z_{t-1}|z_{t},z_{0})$ can be modeled as a Gaussian $\mathcal{N}(z_{t-1};\mu_{t}(z_{t},z_{0}),\sigma_{t}^{2}\mathbf{I})$. The mean of this distribution, $\mu_{t}(z_{t},z_{0})$, can be parameterized as:
@@
By approximating $p_{\theta}(z_{t-1}|z_{t})$ with a Gaussian distribution $\mathcal{N}(z_{t-1};\mu_{\theta}(z_{t},t),\sigma_{t}^{2}\mathbf{I})$, the KL-divergence is simplified as minimizing the difference between $\mu_{t}(z_{t},z_{0})$ and $\mu_{\theta}(z_{t},t)$. Parameterizing $\mu_{\theta}(z_{t},t)$ as
$\mu_{\theta}(z_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}(z_{t}-\epsilon_{\theta}(z_{t},t)\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}),$ (13)
the problem is further simplified as accurately approximating the noise component $\epsilon$, which is facilitated by a neural network denoted as $\epsilon_{\theta}$. The network is typically conditioned on a context variable $c$ that can embed additional information such as text descriptions. Consequently, the network $\epsilon_{\theta}$ is trained to minimize the discrepancy between the true noise $\epsilon$ and its estimated value, using the loss function:
@@ II-C2 The Denoiser Structure
The structure of the proposed denoiser, $\epsilon_{\theta}$, is illustrated in Figure 3. The UNet architecture [47] of the Stable Diffusion denoiser is depicted in blue and remains unchanged during training. This architecture comprises of four encoder blocks, four decoder blocks, and one middle block. Each encoder/decoder block includes two main blocks and one convolutional block for scaling operations. The middle block contains two main blocks. Each main block consists of 4 ResNet and 2 Vision Transformers (ViTs). The ViTs integrate self-attention and cross-attention mechanisms to facilitate interaction with external prompts, such as text encoded by the CLIP text encoder.
@@
The control module highlighted in green in Figure 3 is a pivotal component of the proposed denoiser architecture. Following the ControlNet’s structure [41], this module adopts the UNet encoder and middle block in the Stable Diffusion denoiser and is initialized with the pre-trained weights. The control module is comprised of four encoder blocks in addition to the middle block to handle four distinct resolutions—64×$\times 64$ 64, $32\times 32$, $16\times 16$, and $8\times 8$—capturing a wide range of feature abstraction scales. The outputs from these blocks are fed into the four decoder blocks and the middle block of the Stable Diffusion UNet. Before being fed into the UNet, the control module’s outputs go through a series of $1\times 1$ convolutional layers (‘zero conv’ layers in Figure 3).
@@
It is worth noting that the pre-trained Stable Diffusion UNet (in blue in Figure 3) in the denoiser is adopted without retraining whereas the control module (in green) is trained for the proposed control module input. We obtain the control module input by concatenating the latent representation $z_{t}$ with the spatial features $f_{v}$ from the LDM encoder $E_{ldm}$ as in [39]. This concatenation enhances the input’s representational capacity but it extends the control module network’s channel depth necessitating additional parameters beyond those present in the pre-trained model. To handle this extension, the extra parameters are set to zero at initialization to maintain a neutral effect at the start of the training process.
To encode the channel CSI, we first perform a sinusoidal encoding to the channel SNR $\gamma$ under the assumption that it lies within a fixed range. Then we concatenate it with the real and imaginary part of the channel gain $h$ (characteristic of a Rayleigh Fading channel) before processing through a multi-layer perception (MLP) to produce the CSI embeddings. These embeddings share the same dimensionality with the time embeddings. In the control module, we integrate the CSI embeddings with the time embeddings by addition to inject the channel information. Furthermore, textual features $f_{t}$ serve as crucial prompts infusing contextual information to both the pre-trained UNet and the control module. As shown in Figure 3, textual features $f_{t}$ are used as text prompts and are passed to the cross-attention layers of ViTs within each encoder/middle/decoder block as keys and values.
Similar to the LDM loss (14), our approach focuses on reducing the Mean Square Error (MSE) of the noise prediction while adhering to the condition set $\bar{c}$ given in equation (9). Our loss function is thus represented by:
@@
  Input: Time steps $T$, spatial features $f_{v}$, textual features $f_{t}$, channel SNR $\gamma$, channel gain $h$, scale factor $\lambda$, conditional denoiser $\epsilon_{\theta}$
@@
       $\hat{z}_{0}\leftarrow\tilde{z}_{0}-\frac{\lambda}{C_{l}H_{l}W_{l}}(\tilde{z}_{0}-f_{v})$
@@ II-C3 Adding Intermediate Guidance
Though the proposed DiffJSCC does not inherently rely on intermediate guidance throughout the diffusion process, as articulated by Lin et al. [39], diffusion models strive to carefully navigate the delicate trade-off between the realism of the generated images and their fidelity to the original content. It is worth mentioning that these models might unintentionally introduce mismatched elements that may look realistic but different from the actual target. This issue is more prevalent when there is a significant difference between the target and training domains. In scenarios where it is crucial to ensure that reconstructed images faithfully reflect the original content, we incorporate an optional guidance mechanism that leverages the initial JSCC reconstruction $\hat{x}$ to direct the intermediate latent state $\tilde{z}_{0}$ during the denoising process. Prior to applying guidance, the intermediate state $\tilde{z}_{0}$ at a particular time $t$ can be determined equation (11):
@@
$\hat{z}_{0}=\tilde{z}_{0}-\frac{\lambda}{C_{l}H_{l}W_{l}}(\tilde{z}_{0}-f_{v}),$ (17)
where the variables $C_{l}$, $H_{l}$, and $W_{l}$ correspond to the dimensions of the latent space, and $\lambda$ is a scalar that regulates the level of alignment between $\hat{z}_{0}$ and $f_{v}$. The posterior distribution $q(z_{t-1}|z_{t},\hat{z}_{0})$ subsequently determines the sampled state $z_{t-1}$ as given by equation (12) substituting $z_{0}$ with $\hat{z}_{0}$. This entire process is captured in Algorithm 1. It is worth emphasizing that the application of the intermediate guidance is optional and can be omitted by assigning a value of zero to $\lambda$, thereby maintaining the general functionality of the diffusion process without the additional guidance.
From the denoised $z_{0}$ (or $\hat{z}_{0}$), the final output image $\hat{x}_{diff}$ is generated using an LDM decoder $\hat{x}_{diff}=D_{ldm}(z_{0})$.
@@ III Training and Evaluation
Figure 4: Evaluation of the proposed approach versus baselines across different transmission rates under 1dB (top) and 10dB SNR (bottom) on the Kodak dataset under AWGN channel. Lower LPIPS/FID scores indicate higher perceptual quality.
Figure 5: Evaluation of the proposed approach versus baselines across different SNR levels with transmission rates $\rho$ of $1/384$ (top) and $1/96$ (bottom) using the Kodak dataset under the AWGN channel. Lower LPIPS/FID scores indicate higher perceptual quality.
Figure 6: Evaluation of the proposed approach versus baselines across different SNR levels with transmission rates $\rho$ of $1/192$ (top) and $1/48$ (bottom) using the Kodak dataset under the Rayleigh Fading channel. Lower LPIPS/FID scores indicate higher perceptual quality.
Figure 7: Visual comparison between DiffJSCC and baselines on the Kodak dataset at 1dB SNR under the AWGN channel. For DiffJSCC, we show both the final reconstructed images and the captions generated from the initial reconstruction. The keywords are marked in red. Zoom in for better perception.
@@ III-A1 Dataset
Our DiffJSCC framework evaluation encompasses two principal applications: general image transmission and specialized facial image transmission. In the general image transmission scenario, DiffJSCC is fine-tuned on a selected dataset of approximately 300,000 high-resolution images from the OpenImages collection [^1] https://storage.googleapis.com/openimages/web/index.html. Throughout the training phase, these images are subject to random cropping operations, resulting in dimensions of $256\times 256$ pixels. For the performance evaluation, we use the standard Kodak dataset, which consists of 24 curated high-fidelity images with dimensions of $768\times 512$ pixels. In addition, the performance evaluation is extended to the ADE20K validation set with 2000 images to assess image reconstruction quality and semantic segmentation accuracy.
@@ III-A2 Implementation
Within the proposed framework, we utilize an SNR-adaptive JSCC architecture for both the encoder and decoder, as detailed by Yang et al.[23]. This architecture incorporates the CSI into the feature processing pipeline via a Squeeze-Excitation mechanism [22], enhancing the adaptability to varying SNR conditions. The transmission rate $\rho$ is fixed for each model configuration without the policy network introduced in [23]. Within this network structure, $\rho$ is determined by the output channel count ($C_{out}$) and the downsampling factor ($D$), satisfying $\rho=C_{out}/(3\times 2^{2D+1})$. General image transmission training involves random $256\times 256$ patches and a batch size of 32, while the facial image dataset utilizes full $512\times 512$ CelebAHQ images with a smaller batch size of 16. An Adam optimizer initializes learning at $10^{-3}$, tapering over 100,000 iterations following a cosine decay schedule.
@@
We used 2 NVIDIA A40 GPUs to accelerate training. Though the Stable Diffusion model is trained using 1000 time steps, we use a spaced DDPM sampling for 50 steps during inference adopting a common strategy [43, 39] to improve sampling efficiency. Unless specified otherwise, the scaling factor $\lambda$ is preset to zero as a standard baseline configuration without the intermediate guidance introduced in Section II-C3.
@@ III-B Metrics
For evaluations, we adopt standard distortion metrics such as PSNR and MS-SSIM[48] to measure image fidelity concerning pixel intensity and structural details, respectively. To evaluate the perceptual quality of reconstructed images, we employ more recently proposed metrics such as Learned Perceptual Image Patch Similarity (LPIPS)[49] and Fréchet Inception Distance (FID) scores[50]. These metrics provide crucial insights into visual similarity and how well the distribution of generated images aligns with that of real images. Given the Kodak dataset contains only 24 images, we transmit each image five times to measure our performance metrics reliably. For Kodak images, the FID score is calculated over half-overlapping patches of $256\times 256$ pixels. Our evaluation also includes the performance of the semantic segmentation task on the generated and original images measured by the mean Intersection over Union (mIoU) metric.
@@ III-C Performance on General Image Transmission
DeepJSCC and DiffJSCC methods support four distinct rates: $\rho=1/384,1/192,1/96,$ and $1/48$. Notably, the basic DeepJSCC method surpasses traditional separate coding methods across multiple metrics, validating previous studies. Though the diffusion-based refinements from our DiffJSCC may induce degradation in traditional distortion metrics such as PSNR and MS-SSIM similar to [30, 39], DiffJSCC exhibits remarkable improvements in visual quality assessed by LPIPS and FID metrics. These metrics capture the capability of DiffJSCC to produce better photorealistic reconstructions even at compression levels unachievable by the other schemes (including theoretical ‘BPG-Capacity’).
@@
Notice, at a 1dB SNR, DiffJSCC attains a low FID score of $117.9$ at the transmission rate of merely $\rho=1/384=0.0026$, outperforming the ‘BPG+Capacity’ baseline’s FID score of $165.4$ at a substantially higher rate of $\rho=0.0377$. This translates to DiffJSCC achieving superior image quality by a margin of $28.7\%$ while utilizing 14×14$\times$14 × less symbols. In scenarios where the SNR is increased to 10dB, DiffJSCC continues to hold a significant advantage over the other methods although the performance gap in terms of LPIPS and FID is reduced. When transmitting at a rate of $\rho=0.0026$, the FID score of DiffJSCC demonstrates an impressive $64.3\%$ enhancement over DeepJSCC and a $70.7\%$ over ‘BPG+Capacity’. This confirms the effectiveness of the DiffJSCC framework in reconstructing highly realistic images under stringent rate and channel constraints.
@@
In Figure 5 and 6, we assess DiffJSCC in comparison to other baseline approaches across various SNR conditions for AWGN channel and Rayleigh Fading channel respectively. The DeepJSCC-$adv$ model from Yang et al.[32] is included for comparison, which uses the hyperparameter $\alpha$ to adjust the weight of adversarial loss. Various $\alpha$ parameters were tested to explore the trade-off between distortion and visual quality. Similarly, we experimented with different $\lambda$ values in DiffJSCC with the same goal. These evaluations were conducted at transmission rates of $\rho=1/384$ and $\rho=1/96$, with SNR values set at $\gamma=1,4,7,10,13$ dB. All benchmarks including ours are based on the same deep JSCC network architecture. The results in Figure 5 demonstrate that although the DeepJSCC-$adv$ baseline successfully lowers LPIPS, it struggles to effectively minimize FID scores regardless of the $\alpha$ value used. The FID score remained above $190$ at $\rho=1/384$ and exceeded $110$ at $\rho=1/96$. Conversely, DiffJSCC significantly improves both LPIPS and FID scores over the entire range of SNR values, while maintaining similar PSNR and MS-SSIM distortion metrics. At $\lambda=0$ and $\rho=1/384$, DiffJSCC shows a notable average improvement of $41.4\%$ in LPIPS and $56.7\%$ in FID scores compared to the standard DeepJSCC. When compared with DeepJSCC-$adv$ using $\alpha=5\times 10^{-3}$, it achieves $21.1\%$ better LPIPS and $52.2\%$ enhanced FID metrics. At $\rho=1/96$, DiffJSCC consistently demonstrates a significant gain in performance, improving average LPIPS and FID scores by $43\%$ and $53.2\%$ over DeepJSCC, respectively, and $18.9\%$ and $48.4\%$ compared to DeepJSCC-$adv$. Increasing the scale factor $\lambda$ from 0 to 100 gradually aligns the generated image with the initial JSCC reconstruction, enhancing PSNR and MS-SSIM but reducing perceptual quality. However, this trade-off still leads to better visual perception metrics (LPIPS and FID) than the baselines. This significant improvement in the perceived visual quality highlights the effectiveness of our fine-tuning control module along with the Stable Diffusion model. A similar trend can be observed under the Rayleigh Fading channel as shown in Figure 6.
@@
Figure 7 provides visual examples of the superior perceptual quality achieved by DiffJSCC at a challenging SNR of 1dB under the AWGN channel. At a transmission rate of $\rho=1/384$, the reconstructions produced by the conventional DeepJSCC framework exhibit significant distortion and have difficulty recovering important image details. Conversely, DiffJSCC demonstrates a remarkable capacity to restore crucial elements of the image, like the structures of the lighthouse and house, offering significantly enhanced perceptual quality despite slight variances from the original. With the transmission rate raised to $1/96$, DiffJSCC’s capabilities become more evident, resulting in reconstructions that accurately reflect the original images. Furthermore, we demonstrate the validity of captions generated by the BLIPv2 model based on the initial JSCC reconstructions. These captions prove to be robust against variations in reconstruction quality, consistently capturing the important content keywords of each image (highlighted in red). This confirms the ability of DiffJSCC to convey essential semantic information even at reduced transmission rates.
Figure 8: Results on the ADE20k dataset with $\rho=1/384$ and $\rho=1/96$. Higher mIoU scores indicate better segmentation.
Figure 9: Visualizations on ADK20K with $\rho=1/384$ and 1dB SNR under the AWGN channel.
@@
Figure 10: Comparison between the proposed method and baselines at various SNRs on CelebAHQ under the AWGN channel.
@@
Figure 11: Visualizations on CelebAHQ with $\rho=1/768$ in both high and low SNR scenarios under the AWGN channel.